# Generalized Brillouin Zone Fragmentation

### v1

@@ Abstract
where $z=e^{ik}$, $k$ the lattice momentum. For $b$ atoms per unit cell, $H_{j}$ is a $b\times b$ matrix representing the hopping amplitudes between atoms $j$ unit cells apart, with maximal left and right hopping ranges $q$ and $p$. In a finite $N$-unit cell setting, $H(z)$ is represented in real space by a $bN\times bN$ block Toeplitz matrix with $j$-th diagonals given by $H_{j}$; for periodic boundary conditions (PBCs), $j$ is taken modulo $N$.
@@
and $\varphi_{\mu}$ is its corresponding $b$-component intra unit-cell eigenstate satisfying $H(z_{\mu})\varphi_{\mu}=E\varphi_{\mu}$. The offset $x_{0}$ controls the position at which the relative bulk weights $c_{\mu}$ are evaluated: Setting $x_{0}=0$ or $N$ reveals the bulk eigensolution composition at the edges, while setting $x_{0}=N/2$ yields a symmetrically-defined bulk composition.
Importantly, since each $\phi_{\mu}(x)$ is already a bulk eigensolution, the OBC eigenvalue problem is reduced to that of determining coefficients $c_{\mu}$ that satisfy the constraints from boundary hopping truncations, a problem that does scale with $N$. As detailed in Sect. I of [76], these constraints can be succinctly written as $\sum_{\mu}M_{\nu\mu}c_{\mu}=0$, where
@@
However, beyond the simplest models, the kernel of $M$ can be much more intricate, even depending strongly on $N$ [76]. Physically, that corresponds to competing skin accumulations within the lattice that can non-trivially distort the non-Bloch picture. When the non-vanishing $c_{\mu}$ coefficients correspond to more than one inverse skin decay rate $\log|z_{\mu}|$, we say that the GBZ has fragmented – this is vividly illustrated in Fig. 1, where the spectra (Top Row) of illustrative models break up into multiple constituent $|z|$ “layers” (2nd Row). To quantify the extent of GBZ fragmentation at each $E_{\text{OBC}}$, we introduce the composition-IPR
@@
where $1/\text{cIPR}$ measures the effective number of fragmented GBZ pieces. For conventional GBZs, the cIPR typically lies between $0.5$ and $1$ (it is $0.5$ when the two nonzero coefficients $c_{\mu}$ are equally large).
Occurrence of fragmented GBZs.– As a warm-up, we first examine the simplest Hatano-Nelson (HN) model $H_{\text{HN}}(z)=z+h/z$, which does not exhibit GBZ fragmentation. Its left/right open boundary constraints can be succinctly encoded as $M\mathbb{c}=\mathbb{0}$, with $M$ being the $2\times 2$ matrix $\left(\begin{matrix}h/z_{1}^{x_{0}+1}&h/z_{2}^{x_{0}+1}\\ z_{1}^{N-x_{0}}&z_{2}^{N-x_{0}}\end{matrix}\right)$ in the bulk solution basis [Eq. 2[76]. Simultaneously satisfying both boundary conditions requires the basis coefficients $\mathbb{c}=(c_{1},c_{2})^{T}$ to satisfy $c_{1}/c_{2}=-\left(z_{2}/z_{1}\right)^{N-x_{0}}$. But DetM=0M=0 necessitates $z_{1}^{N+1}=z_{2}^{N+1}$, which fixes $|c_{1}/c_{2}|=1$ for all $N$. Furthermore, as $E_{\text{OBC}}=z+h/z$ is symmetric under $z\leftrightarrow h/z$, the GBZ is determined to be $z_{1}=z_{2}^{*}=\sqrt{h}e^{ik}$, a classic example of a non-fragmented GBZ  [^1] This also holds for the non-Hermitian SSH model, which has the same form of characteristic dispersion polynomial..
@@
Figure 4: (a) GBZ fragmentation occurs in simple metamaterial structures with multiple eigenmodes per unit cell, as in our illustrative photonic crystals (see [76] for details). (b) Highly complicated diffuse unit cell eigenmodes lead to multi-component heavily-coupled effective lattices reminiscent of $H_{\text{rand}}$. (c) The relative entropy $S$, which measures the contrast between PBC/OBC spectra, is much lower for anti-aligned cases (dashed) at small $g$, and also universally decreases with system size $N$. It is far lower than that of $H_{\text{HN}}$ (black), which experiences clean GBZ deformations.
We consider PhCs under aligned/anti-aligned configurations [Fig. 4(a)], which reduce to 1D two-chain models with parallel and competing NHSE, coupled more strongly for smaller transverse spacing $g$. As shown in Sect. V of [76], their eigenfrequency spectra exhibits stark resemblance to that of $H_{\text{rand}}$ [Fig. 1(d)], with complicated random-looking structures and characteristic dependence on $N$. To quantify the NHSE extent in such intricate GBZ-fragmented spectra, we convert the PBC/OBC spectra to distributions $p_{\text{PBC/OBC}}(\omega)$, and compute their symmetrized relative entropy [83] $S=\int\left(p_{\text{PBC}}(\omega)-p_{\text{OBC}}(\omega)\right)\log\frac{p_{\text{PBC}}(\omega)}{p_{\text{OBC}}(\omega)}d\omega$ [Sect. IV of [76]]. Indeed, from Fig. 4(c), increasing $N$ universally leads to lower $S$ (suppressed NHSE), as expected from stronger NHSE competition under both configurations. At smaller $g$ where inter-chain coupling dominates, anti-aligned cases (dashed) also exhibits much lower $S$, consistent with GBZ fragmented systems [76].
@@
Acknowledgements.– HYM acknowledges support by the Natural Science Foundation of Hunan Province (2023JJ40612). CHL acknowledges support from the Ministry of Education, Singapore (MOE Tier-II award number: MOE-T2EP50222-0003) and (Tier-I WBS number: A-8002656-00-00). YSA acknowledges support from the Ministry of Education, Singapore, under the award number MOE-T2EP50224-0021.
@@
Conventionally, to solve for the OBC spectrum and eigenstates in a finite OBC setting with $N$ unit cells, one would need to diagonalize the bounded $bN\times bN$ block Toeplitz matrix $H_{\text{OBC}}$ whose $(x,x^{\prime})$-th block is given by $H_{x-x^{\prime}}$. This, however, is computationally expensive since the computation time typically scales polynomially [92] with $bN$, which can be substantial if $N$ is large. Furthermore, in the presence of the non-Hermitian skin effect (NHSE), floating point errors are also amplified exponentially with $N$.
@@
This drastic reduction hinges on the fact that, in the bulk eigenstate basis, only $B$ boundary constraints need to be explicitly satisfied, even as the system size $N$ becomes arbitrarily large. Given any reference energy $E$, one can identify bulk eigenstates $\phi_{\mu}(x)$ that generically assume the form
@@
To summarize, when $E$ is an OBC eigenenergy possessing an eigenstate $\psi_{\text{OBC}}(x)=\sum_{\mu}c_{\mu}z_{\mu}^{x}\varphi_{\mu}$, the coefficients $c_{\mu}$ lies in the kernel of the $M$, which necessitate that $\text{det}M=0$. Hence, to check whether a given energy $E$ belongs to the OBC spectrum of an arbitrary large $N$ unit-cell system, one just needs to obtain the (non-Bloch) bulk eigenstates $\varphi_{\mu}$ and their $z_{\mu}$ [Eq. S3], and use them to compute the $B\times B$ matrix $M$ [Eq. S5]. The point $E$ belongs to the OBC spectrum if $\text{Det}M$ vanishes, as illustrated in Fig. S1 (in practice, it just needs to be a sharp local minimum, since departures from $\text{Det}M=0$ (including floating point errors) generally scale with the power of $N$).
@@
where $c_{\mu}$ is the weightage of the bulk $\phi_{\mu}$ eigensolution in the corresponding $\psi_{\text{OBC}}$, and can be obtained by solving $M\mathbb{c}=\mathbb{0}$ with $M$ given by Eq. S5. When only two $z_{\mu}$ solutions have non-negligible weights $|c_{\mu}|=\cos\theta$ and $|c_{\nu}|=\sin\theta$, we have $\text{cIPR}=\cos^{4}\theta+\sin^{4}\theta=(3+\cos 4\theta)/4$, which lies between $0.5$ and $1$. In (hypothetical) other extreme where each coefficient is equally weighted, we have $|c_{\mu}|=1/\sqrt{B}$ and $\text{cIPR}=B^{-1}\cong 1/b(p+q)$. In the various examples shown in this work, we typically have cIPR between 0.2 and 0.8.
@@
    To further derive the exact (slight) dependence of $z_{0}$ and $\theta$ on the system size $N$, we invoke
@@
    Note that Eq. S15 remains the same even if an additional NN hopping is introduced to give nonvanishing $H_{1}$. To proceed, we note that the middle term does not contain any power of $z_{1}^{N}$, and will thus be dominated by the other two terms at large $N$. Hence, we obtain $z_{1}^{N+1}z_{2}^{N+1}(z_{2}-z_{1})=z_{1}^{N+1}z_{3}^{N+1}(z_{3}-z_{1})$, which yields
@@
    with index $n$ running within the range where $|\theta|<\frac{\pi}{3}$. Interestingly, the denominator is not $N$ as in usual PBC Bloch waves, or $N+1$ as in the OBC Hatano-Nelson model, but $N+\frac{4}{3}$, which includes the shift meticulous obtained from the transcendental Eq. S17. In this sense, it appears that $\theta$ can be regarded as a “momentum” index corresponding to the wavenumbers of non-Bloch states in a slightly “expanded” space. These states can still be consistent with the boundary conditions due to the presence of the small $c_{1}$ coefficient.
@@
    where $\eta=1$ or $e^{\pm 2\pi i/3}$ is a cube root of unitary, and $\theta$ is given by Eq. S18. Its accuracy is demonstrated in Fig. S3(a): The numerical eigenvalues (dark blue) exhibit excellent agreement with analytically computed eigenvalues from Eq. S19, with $\theta$ given by Eq. S18 (light cyan). However, if $\theta$ is computed without replacing $N$ by $N+4/3$, then the computed eigenvalues (green) no longer coincide.
@@
    At very small $N$, the system essentially behaves like two opposite uncoupled HN chains, with a total of 4 constant GBZ solutions, $|z_{+}|=|z_{-}|$ and $|z_{+}|^{-1}=|z_{-}|^{-1}$. However, beyond a critical system size $N$, the GBZ solutions bifurcate (fragmentate) into non-degenerate solutions $z_{+},z_{-},1/z_{+},1/z_{-}$, as shown in Fig.  S4(a). In general, these solutions then converge very slowly to their thermodynamic limit values, remaining fully fragmented unless $|z_{-}|$ hits unity (as in the illustrative example in Fig. S4(a)). Below, we derive these behavior in detail.
    To compute $M$, Eq. S5 gives $M_{1\mu}=z^{-1-x_{0}}_{\mu}H_{-1}\varphi_{\mu}$ and $M_{2\mu}=z^{N-x_{0}}_{\mu}H_{1}\varphi_{\mu}$ where $N$ is the system size, which is similar in form to the previous case since we also have $p=q=1$. However, due to the very different $\varphi_{\mu}$ eigenvectors, $M$ would turn out very different. Writing $\varphi_{\mu}=\left(\begin{matrix}\varphi_{\mu}^{\uparrow}\\ \varphi_{\mu}^{\downarrow}\end{matrix}\right)$, $M\mathbb{c}=\mathbb{0}$ is equivalently expressed as
@@
    where $z_{\pm}$ are the two roots adiabatically connected to the two solutions of $\left(z+\frac{h}{z}-E_{\text{OBC}}\right)$ in Eq. S29. Together, Eq. S29 and Eqs. S26 or S27 constitute two constraints that allow the $E_{\text{OBC}}$ spectrum to be solved in terms of $\Delta$ and $N$ (typically numerically). Specifically, for any provisional $E_{\text{OBC}}$ on the 2D complex plane, one can first find its $z_{\pm}$ solutions using Eq. S29. Then Eqs. S26 or S27 can be applied to check whether such a solution is indeed valid. This constraints the valid $E_{\text{OBC}}$ solutions to a co-dimension one subspace, which is just the set of OBC spectral curves.
    Figure S4: GBZ fragmentation in the coupled-HN model [Eq. S21], all with $h=0.6$. (a,b) While the GBZ is constant and doubly degenerate (non-fragmented) at small $N$, it bifurcates and becomes fragmented (with $|z_{+}|\neq|z_{-}|$ ) beyond that. For the $\Delta=0.05$ coupling used, the $z_{-}$ solution subsequently becomes constant and doubly degenerate again after it meets with $1/z_{-}$. (b) The composition ratio $R=c_{+}/c_{-}$ for the $z_{+}$ vs. $z_{-}$ solutions obeys an asymptotic exponential relationship with $N$, as theoretically given (black curve) in Eq. S25 and its accompanying table, and numerically verified (black dots) here for the largest Re(E)(E) eigenstate with $\Delta=0.05$. (c) Accurate prediction of the fragmentation of $z_{+}$(blue) and $z_{-}$(orange) for the illustrative $E_{\text{OBC}}=-1.512-0.0519i$ eigenstate at $N=30$. The numerical data (dots) closely align with their theoretical predictions (solid curves) $|z_{+}|=1.76\Delta^{1/(N+1)}$, $|z_{-}|=0.947\Delta^{-1/(N+1)}$ from Eq. S31.
@@
    which resembles logarithmic behavior in the limit of large $N$. Very good agreement with numerics is demonstrated in Fig. S4(c) for an illustrative eigenstate with complex eigenvalue. Similar fits can be obtained for other eigenstates with complex energies; for real eigenergies, $|z_{+}|=|z_{-}|=\sqrt{h}$ remains constant, as if the chains are uncoupled.
@@
However, to have higher degeneracies i.e. $|z|=|z^{\prime}|$ and $|z^{\prime}|=|z^{\prime\prime}|$, we require the coincidence of more than one such constraint. In general, that can only exist at the intersections of 1D solution curves, i.e. isolated points. To our knowledge, unlike in the quadratic case, there is no single constraint that can give rise to $3$ or more roots related by multiplication of the respective root of unity [95].
@@
Figure S6: Evolution of the GBZ with system size NN for various models, from $N=5$ unit cells (green) to $N=30$ (pink). More intense colors correspond to greater GBZ weight $|c_{\mu}|^{2}$ computed at $x_{0}=N/2$ [Eqs. S4 and S5]. Shown are models (a-c) $H_{\text{coupled-HN}}$ [Eq.S38] and (d) $H_{\text{coupled-topo}}$ [Eq.S39], all containing subchains with antagonistic NHSE. All GBZs shown exhibit significant evolution with $N$, except in (c) where only the more fragmented (fainter) top and bottom regions shift with $N$.
@@
Figure S7: Variation of the spectra, GBZ composition and GBZ of $H_{\text{coupled-topo}}$ [Eq. S39] across different system sizes $N$, with coupling $\Delta=10^{-4}$, hoppings $t=0.6$, $t^{\prime}=0.1$ and asymmetry $\gamma=2$. Significant spreading i.e. fragmentation (yellow $\rightarrow$ purple) of the bulk contributions across different $|z|$ occur at all NN, not just confined to the topological transition.
@@
Since the $M$ matrix depends on $N$ [Eq. S5], the allowed $E_{\text{OBC}}$ and $c_{\mu}$ composition coefficients also generically depend significantly on the system size $N$, unless the usual GBZ condition (where only two $z_{\mu},z_{\nu}$ solutions dominate) is met.
The system size dependence of the GBZs of $H_{\text{coupled-HN}}$ and $H_{\text{coupled-topo}}$ are shown in Fig. S6, for illustrative parameters. In general, small inter-chain couplings $\Delta$ give rise to stronger variations with $N$ (a,b,d), even though the GBZs still do not converge into one non-fragmented ring at large $N$ (magenta). When the chains are weighted unequally by $g\neq 1$ (b), the GBZ rings also break up into qualitatively distinct segments with different $N$-dependent behavior, a result of the interplay between dissimilar hopping directions and strength. At the special choice of strong coupling $\Delta=0.3$ which prevents any amplification in the translation-invariant setting (such that the PBC spectrum becomes real), the GBZ is still not entirely Bloch (on the unit circle), and some fragmentation and variation with $N$ can be seen at the top and bottom regions.
The evolution of the GBZ with $N$ for $H_{\text{coupled-topo}}$ is further elaborated in Fig. S7, in the weak coupling regime of $\Delta=10^{-4}$. Its spectrum (Top Row) transitions from topologically gapped to a non-topological point gap (loop) as $N$ is increased, with some GBZ fragmentation (magenta) mostly observed near the gap crossing. Along these states, some significant distribution of bulk contributions $\phi_{\mu}$ across different $|z|$ can be observed (Middle Row). Correspondingly, they appear as region of fuzziness (faint yellow) in the GBZ loci (Bottom Row). These are also the same regions that ”absorb” the topological states (isolated dots) as the topological transition is crossed.
@@
Figure S8: GBZ and spectra of illustrative $b$-component models $H_{\text{rand}}$ with randomly-generated hoppings across all atoms within and across $N$ unit cells [Eq. S40], all with $N=30$ unit cells. As $b$ increases, fragmentation (faint blue in the GBZ and non-red eigenvalues in the spectrum) becomes inevitable.
@@
For completeness, the scaling behavior of the GBZ composition for this random $5$-component model is also plotted in Fig. S10. Even at large sizes $N$, some eigenstates maintain their fragmented GBZ nature, appearing blurry in $z$.

### v2

@@ Abstract
where $z=e^{ik}$, $k$ the lattice momentum. For $b$ atoms per unit cell, $H_{j}$ is a $b\times b$ matrix representing the hopping amplitudes between atoms $j$ unit cells apart, with maximal left and right hopping ranges $q$ and pp. In a finite NN-unit cell setting, $H(z)$ is represented in real space by a $bN\times bN$ block Toeplitz matrix with $j$-th diagonals given by $H_{j}$; for periodic boundary conditions (PBCs), $j$ is taken modulo NN.
@@
and $\varphi_{\mu}$ is its corresponding $b$-component intra unit-cell eigenstate satisfying $H(z_{\mu})\varphi_{\mu}=E\varphi_{\mu}$. The offset $x_{0}$ controls the position at which the relative bulk weights $c_{\mu}$ are evaluated: Setting $x_{0}=0$ or NN reveals the bulk eigensolution composition at the edges, while setting $x_{0}=N/2$ yields a symmetrically-defined bulk composition.
Importantly, since each $\phi_{\mu}(x)$ is already a bulk eigensolution, the OBC eigenvalue problem is reduced to that of determining coefficients $c_{\mu}$ that satisfy the constraints from boundary hopping truncations, a problem that does scale with NN. As detailed in Sect. I of [76], these constraints can be succinctly written as $\sum_{\mu}M_{\nu\mu}c_{\mu}=0$, where
@@
However, beyond the simplest models, the kernel of $M$ can be much more intricate, even depending strongly on NN [76]. Physically, that corresponds to competing skin accumulations within the lattice that can non-trivially distort the non-Bloch picture. When the non-vanishing $c_{\mu}$ coefficients correspond to more than one inverse skin decay rate $\log|z_{\mu}|$, we say that the GBZ has fragmented – this is vividly illustrated in Fig. 1, where the spectra (Top Row) of illustrative models break up into multiple constituent $|z|$ “layers” (2nd Row). To quantify the extent of GBZ fragmentation at each $E_{\text{OBC}}$, we introduce the composition-IPR
@@
where $1/\text{cIPR}$ measures the effective number of fragmented GBZ pieces. For conventional GBZs, the cIPR typically lies between $0.5$ and 11 (it is $0.5$ when the two nonzero coefficients $c_{\mu}$ are equally large).
Occurrence of fragmented GBZs.– As a warm-up, we first examine the simplest Hatano-Nelson (HN) model $H_{\text{HN}}(z)=z+h/z$, which does not exhibit GBZ fragmentation. Its left/right open boundary constraints can be succinctly encoded as $M\mathbb{c}=\mathbb{0}$, with $M$ being the $2\times 2$ matrix $\left(\begin{matrix}h/z_{1}^{x_{0}+1}&h/z_{2}^{x_{0}+1}\\ z_{1}^{N-x_{0}}&z_{2}^{N-x_{0}}\end{matrix}\right)$ in the bulk solution basis [Eq. 2[76]. Simultaneously satisfying both boundary conditions requires the basis coefficients $\mathbb{c}=(c_{1},c_{2})^{T}$ to satisfy $c_{1}/c_{2}=-\left(z_{2}/z_{1}\right)^{N-x_{0}}$. But DetM=0M=0 necessitates $z_{1}^{N+1}=z_{2}^{N+1}$, which fixes $|c_{1}/c_{2}|=1$ for all NN. Furthermore, as $E_{\text{OBC}}=z+h/z$ is symmetric under $z\leftrightarrow h/z$, the GBZ is determined to be $z_{1}=z_{2}^{*}=\sqrt{h}e^{ik}$, a classic example of a non-fragmented GBZ  [^1] This also holds for the non-Hermitian SSH model, which has the same form of characteristic dispersion polynomial..
@@
Figure 4: (a) GBZ fragmentation occurs in simple metamaterial structures with multiple eigenmodes per unit cell, as in our illustrative photonic crystals (see [76] for details). (b) Highly complicated diffuse unit cell eigenmodes lead to multi-component heavily-coupled effective lattices reminiscent of $H_{\text{rand}}$. (c) The relative entropy $S$, which measures the contrast between PBC/OBC spectra, is much lower for anti-aligned cases (dashed) at small $g$, and also universally decreases with system size NN. It is far lower than that of $H_{\text{HN}}$ (black), which experiences clean GBZ deformations.
We consider PhCs under aligned/anti-aligned configurations [Fig. 4(a)], which reduce to 1D two-chain models with parallel and competing NHSE, coupled more strongly for smaller transverse spacing $g$. As shown in Sect. V of [76], their eigenfrequency spectra exhibits stark resemblance to that of $H_{\text{rand}}$ [Fig. 1(d)], with complicated random-looking structures and characteristic dependence on NN. To quantify the NHSE extent in such intricate GBZ-fragmented spectra, we convert the PBC/OBC spectra to distributions $p_{\text{PBC/OBC}}(\omega)$, and compute their symmetrized relative entropy [83] $S=\int\left(p_{\text{PBC}}(\omega)-p_{\text{OBC}}(\omega)\right)\log\frac{p_{\text{PBC}}(\omega)}{p_{\text{OBC}}(\omega)}d\omega$ [Sect. IV of [76]]. Indeed, from Fig. 4(c), increasing NN universally leads to lower $S$ (suppressed NHSE), as expected from stronger NHSE competition under both configurations. At smaller $g$ where inter-chain coupling dominates, anti-aligned cases (dashed) also exhibits much lower $S$, consistent with GBZ fragmented systems [76].
@@
Acknowledgements.– HYM acknowledges support by the Natural Science Foundation of Hunan Province (2023JJ40612). CHL acknowledges support from the Ministry of Education, Singapore (MOE Tier-II award number: MOE-T2EP50222-0003) and (Tier-I WBS number: A-8002656-00-$0$). YSA acknowledges support from the Ministry of Education, Singapore, under the award number MOE-T2EP50224-0021.
@@
Conventionally, to solve for the OBC spectrum and eigenstates in a finite OBC setting with NN unit cells, one would need to diagonalize the bounded $bN\times bN$ block Toeplitz matrix $H_{\text{OBC}}$ whose $(x,x^{\prime})$-th block is given by $H_{x-x^{\prime}}$. This, however, is computationally expensive since the computation time typically scales polynomially [92] with bNbN, which can be substantial if NN is large. Furthermore, in the presence of the non-Hermitian skin effect (NHSE), floating point errors are also amplified exponentially with NN.
@@
This drastic reduction hinges on the fact that, in the bulk eigenstate basis, only $B$ boundary constraints need to be explicitly satisfied, even as the system size NN becomes arbitrarily large. Given any reference energy $E$, one can identify bulk eigenstates $\phi_{\mu}(x)$ that generically assume the form
@@
To summarize, when $E$ is an OBC eigenenergy possessing an eigenstate $\psi_{\text{OBC}}(x)=\sum_{\mu}c_{\mu}z_{\mu}^{x}\varphi_{\mu}$, the coefficients $c_{\mu}$ lies in the kernel of the $M$, which necessitate that $\text{det}M=0$. Hence, to check whether a given energy $E$ belongs to the OBC spectrum of an arbitrary large NN unit-cell system, one just needs to obtain the (non-Bloch) bulk eigenstates $\varphi_{\mu}$ and their $z_{\mu}$ [Eq. S3], and use them to compute the $B\times B$ matrix $M$ [Eq. S5]. The point $E$ belongs to the OBC spectrum if $\text{Det}M$ vanishes, as illustrated in Fig. S1 (in practice, it just needs to be a sharp local minimum, since departures from $\text{Det}M=0$ (including floating point errors) generally scale with the power of NN).
@@
where $c_{\mu}$ is the weightage of the bulk $\phi_{\mu}$ eigensolution in the corresponding $\psi_{\text{OBC}}$, and can be obtained by solving $M\mathbb{c}=\mathbb{0}$ with $M$ given by Eq. S5. When only two $z_{\mu}$ solutions have non-negligible weights $|c_{\mu}|=\cos\theta$ and $|c_{\nu}|=\sin\theta$, we have $\text{cIPR}=\cos^{4}\theta+\sin^{4}\theta=(3+\cos 4\theta)/4$, which lies between $0.5$ and 11. In (hypothetical) other extreme where each coefficient is equally weighted, we have $|c_{\mu}|=1/\sqrt{B}$ and $\text{cIPR}=B^{-1}\cong 1/b(p+q)$. In the various examples shown in this work, we typically have cIPR between 0.2 and 0.8.
@@
    To further derive the exact (slight) dependence of $z_{0}$ and $\theta$ on the system size NN, we invoke
@@
    Note that Eq. S15 remains the same even if an additional NN hopping is introduced to give nonvanishing $H_{1}$. To proceed, we note that the middle term does not contain any power of $z_{1}^{N}$, and will thus be dominated by the other two terms at large NN. Hence, we obtain $z_{1}^{N+1}z_{2}^{N+1}(z_{2}-z_{1})=z_{1}^{N+1}z_{3}^{N+1}(z_{3}-z_{1})$, which yields
@@
    with index $n$ running within the range where $|\theta|<\frac{\pi}{3}$. Interestingly, the denominator is not NN as in usual PBC Bloch waves, or $N+1$ as in the OBC Hatano-Nelson model, but $N+\frac{4}{3}$, which includes the shift meticulous obtained from the transcendental Eq. S17. In this sense, it appears that $\theta$ can be regarded as a “momentum” index corresponding to the wavenumbers of non-Bloch states in a slightly “expanded” space. These states can still be consistent with the boundary conditions due to the presence of the small $c_{1}$ coefficient.
@@
    where $\eta=1$ or $e^{\pm 2\pi i/3}$ is a cube root of unitary, and $\theta$ is given by Eq. S18. Its accuracy is demonstrated in Fig. S3(a): The numerical eigenvalues (dark blue) exhibit excellent agreement with analytically computed eigenvalues from Eq. S19, with $\theta$ given by Eq. S18 (light cyan). However, if $\theta$ is computed without replacing NN by $N+4/3$, then the computed eigenvalues (green) no longer coincide.
@@
    At very small NN, the system essentially behaves like two opposite uncoupled HN chains, with a total of 4 constant GBZ solutions, $|z_{+}|=|z_{-}|$ and $|z_{+}|^{-1}=|z_{-}|^{-1}$. However, beyond a critical system size NN, the GBZ solutions bifurcate (fragmentate) into non-degenerate solutions $z_{+},z_{-},1/z_{+},1/z_{-}$, as shown in Fig.  S4(a). In general, these solutions then converge very slowly to their thermodynamic limit values, remaining fully fragmented unless $|z_{-}|$ hits unity (as in the illustrative example in Fig. S4(a)). Below, we derive these behavior in detail.
    To compute $M$, Eq. S5 gives $M_{1\mu}=z^{-1-x_{0}}_{\mu}H_{-1}\varphi_{\mu}$ and $M_{2\mu}=z^{N-x_{0}}_{\mu}H_{1}\varphi_{\mu}$ where NN is the system size, which is similar in form to the previous case since we also have $p=q=1$. However, due to the very different $\varphi_{\mu}$ eigenvectors, $M$ would turn out very different. Writing $\varphi_{\mu}=\left(\begin{matrix}\varphi_{\mu}^{\uparrow}\\ \varphi_{\mu}^{\downarrow}\end{matrix}\right)$, $M\mathbb{c}=\mathbb{0}$ is equivalently expressed as
@@
    where $z_{\pm}$ are the two roots adiabatically connected to the two solutions of $\left(z+\frac{h}{z}-E_{\text{OBC}}\right)$ in Eq. S29. Together, Eq. S29 and Eqs. S26 or S27 constitute two constraints that allow the $E_{\text{OBC}}$ spectrum to be solved in terms of $\Delta$ and NN (typically numerically). Specifically, for any provisional $E_{\text{OBC}}$ on the 2D complex plane, one can first find its $z_{\pm}$ solutions using Eq. S29. Then Eqs. S26 or S27 can be applied to check whether such a solution is indeed valid. This constraints the valid $E_{\text{OBC}}$ solutions to a co-dimension one subspace, which is just the set of OBC spectral curves.
    Figure S4: GBZ fragmentation in the coupled-HN model [Eq. S21], all with $h=0.6$. (a,b) While the GBZ is constant and doubly degenerate (non-fragmented) at small NN, it bifurcates and becomes fragmented (with $|z_{+}|\neq|z_{-}|$ ) beyond that. For the $\Delta=0.05$ coupling used, the $z_{-}$ solution subsequently becomes constant and doubly degenerate again after it meets with $1/z_{-}$. (b) The composition ratio $R=c_{+}/c_{-}$ for the $z_{+}$ vs. $z_{-}$ solutions obeys an asymptotic exponential relationship with NN, as theoretically given (black curve) in Eq. S25 and its accompanying table, and numerically verified (black dots) here for the largest Re(E)(E) eigenstate with $\Delta=0.05$. (c) Accurate prediction of the fragmentation of $z_{+}$(blue) and $z_{-}$(orange) for the illustrative $E_{\text{OBC}}=-1.512-0.0519i$ eigenstate at $N=30$. The numerical data (dots) closely align with their theoretical predictions (solid curves) $|z_{+}|=1.76\Delta^{1/(N+1)}$, $|z_{-}|=0.947\Delta^{-1/(N+1)}$ from Eq. S31.
@@
    which resembles logarithmic behavior in the limit of large NN. Very good agreement with numerics is demonstrated in Fig. S4(c) for an illustrative eigenstate with complex eigenvalue. Similar fits can be obtained for other eigenstates with complex energies; for real eigenergies, $|z_{+}|=|z_{-}|=\sqrt{h}$ remains constant, as if the chains are uncoupled.
@@
However, to have higher degeneracies i.e. $|z|=|z^{\prime}|$ and $|z^{\prime}|=|z^{\prime\prime}|$, we require the coincidence of more than one such constraint. In general, that can only exist at the intersections of 1D solution curves, i.e. isolated points. To our knowledge, unlike in the quadratic case, there is no single constraint that can give rise to 33 or more roots related by multiplication of the respective root of unity [95].
@@
Figure S6: Evolution of the GBZ with system size NN for various models, from $N=5$ unit cells (green) to $N=30$ (pink). More intense colors correspond to greater GBZ weight $|c_{\mu}|^{2}$ computed at $x_{0}=N/2$ [Eqs. S4 and S5]. Shown are models (a-c) $H_{\text{coupled-HN}}$ [Eq.S38] and (d) $H_{\text{coupled-topo}}$ [Eq.S39], all containing subchains with antagonistic NHSE. All GBZs shown exhibit significant evolution with NN, except in (c) where only the more fragmented (fainter) top and bottom regions shift with NN.
@@
Figure S7: Variation of the spectra, GBZ composition and GBZ of $H_{\text{coupled-topo}}$ [Eq. S39] across different system sizes NN, with coupling $\Delta=10^{-4}$, hoppings $t=0.6$, $t^{\prime}=0.1$ and asymmetry $\gamma=2$. Significant spreading i.e. fragmentation (yellow $\rightarrow$ purple) of the bulk contributions across different $|z|$ occur at all NN, not just confined to the topological transition.
@@
Since the $M$ matrix depends on NN [Eq. S5], the allowed $E_{\text{OBC}}$ and $c_{\mu}$ composition coefficients also generically depend significantly on the system size NN, unless the usual GBZ condition (where only two $z_{\mu},z_{\nu}$ solutions dominate) is met.
The system size dependence of the GBZs of $H_{\text{coupled-HN}}$ and $H_{\text{coupled-topo}}$ are shown in Fig. S6, for illustrative parameters. In general, small inter-chain couplings $\Delta$ give rise to stronger variations with NN (a,b,d), even though the GBZs still do not converge into one non-fragmented ring at large NN (magenta). When the chains are weighted unequally by $g\neq 1$ (b), the GBZ rings also break up into qualitatively distinct segments with different NN-dependent behavior, a result of the interplay between dissimilar hopping directions and strength. At the special choice of strong coupling $\Delta=0.3$ which prevents any amplification in the translation-invariant setting (such that the PBC spectrum becomes real), the GBZ is still not entirely Bloch (on the unit circle), and some fragmentation and variation with NN can be seen at the top and bottom regions.
The evolution of the GBZ with NN for $H_{\text{coupled-topo}}$ is further elaborated in Fig. S7, in the weak coupling regime of $\Delta=10^{-4}$. Its spectrum (Top Row) transitions from topologically gapped to a non-topological point gap (loop) as NN is increased, with some GBZ fragmentation (magenta) mostly observed near the gap crossing. Along these states, some significant distribution of bulk contributions $\phi_{\mu}$ across different $|z|$ can be observed (Middle Row). Correspondingly, they appear as region of fuzziness (faint yellow) in the GBZ loci (Bottom Row). These are also the same regions that ”absorb” the topological states (isolated dots) as the topological transition is crossed.
@@
Figure S8: GBZ and spectra of illustrative $b$-component models $H_{\text{rand}}$ with randomly-generated hoppings across all atoms within and across NN unit cells [Eq. S40], all with $N=30$ unit cells. As $b$ increases, fragmentation (faint blue in the GBZ and non-red eigenvalues in the spectrum) becomes inevitable.
@@
For completeness, the scaling behavior of the GBZ composition for this random 55-component model is also plotted in Fig. S10. Even at large sizes NN, some eigenstates maintain their fragmented GBZ nature, appearing blurry in $z$.